Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4c)
menzione - 22:49 Thursday 18 September 2025 (67741) Print this report
Operator Report - Afternoon shift

ITF found locked at LN3 in SCIENCE mode.
At 15:00 UTC started the planned ISC activity on 8MHz increase and SR tx gain increase carried out by Boldrini and Mantovani.
15:07 UTC - COMMISSIONING mode set. Activity concluded at 17:00 UTC.
17:21 UTC - CALIBRATION mode set.
17:33 - 18:27 UTC - planned LSC injection script long test (300s) - T_PADD = 300

SCIENCE mode set.

At 19:12 UTC ITF unlocked due to a strong earthquake in Russia (fig1). EARTHQUAKE mode set.
NE_ID, PR_ID, SIB2, EIB, SPRB, SNEB, SWEB, SDB2 loops opened. Properly closed. Relocking in progress...

ACS
Under suggestion of Was I contacted Soldani to gave me instructions to act on the CEB cold valve and set it from 20 to 15. Action done at 19:34 UTC (fig2).
SUSP
Meanwhile Palaia recovered the PR vertical position.
DAQ
TelescreenArt VPM process crashed. Properly restarted at 19:40 UTC.

Ego truck passages around CEB (UTC):
14:15 - 14:25

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×